OpenStudy (anonymous):

In last night’s basketball game, Sally scored 9 less than twice what Lucy scored. The sum of their scores is 27. How many points did Sally and Lucy make individually? The heights in inches of three basketball players are 3 consecutive integers. If the sum of twice the 1st, 3 times the 2nd, and the 3rd is 437, what are the three heights. A rectangular floor has a perimeter of 54 feet. If the length is 3 more than the width, what are the dimensions of the floor? Solve for the Given Variable: D= D + pm for M D= RT for T

OpenStudy (anonymous):

also, You are purchasing a CD player at a markdown price of $127.50. If the markdown price was 15% off the original price, how much was the CD player originally?

OpenStudy (anonymous):

1st Q: Let S denote sally's score and L denote Lucy's S=2L-9 S+L=27 solve the equations to get their scores

OpenStudy (anonymous):

let a denote the 1st integer. then other integers wil be b=(a+1) and c= (a+2) then, 2a+3b+c=437 get the variable a, then u get all the integers

OpenStudy (anonymous):